# Implicit Bias of Mirror Flow on Separable Data

### v2

@@ 1.1 Informal statement of the main result
Figure 1: Mirror descent is performed using $3$ different potentials on the same toy $2$d dataset. Left: the losses converge to zero. Center: the iterates converge in direction towards $3$ different vectors, the $3$ lines passing through the origin correspond to the $3$ different separating hyperplanes. Right: these directions are each proportional to $\mathrm{arg\ min\ }\phi_{\infty}({\bar{\beta}})$ under the constraint $\min_{i}y_{i}\langle x_{i},{\bar{\beta}}\rangle\geq 1$ for their respective $\phi_{\infty}$’s, as predicted by our theory (Theorem 1). See Section 5 for more details.
@@ 2 Problem set-up
The first two points of the assumption are common when considering mirror descent (Bauschke et al., 2017) and the third is necessary in continuous-time to ensure the existence and uniqueness over $\mathbb{R}_{\geq 0}$ of a solution to the (MF) differential equation (see Lemma 2 in Appendix A). The coercive gradient assumption is crucial for our main result and we discuss it in more depth in Section 6.
@@ 3 Intuitive construction of the implicit regularisation problem
From now on, we drop the tilde notation and assume that a change of time scale has been done. We want to characterise the directional limit of the diverging iterates $\beta_{t}$. To do so, we study their normalisation $\bar{\beta}_{t}\coloneqq\frac{\beta_{t}}{\|\beta_{t}\|}$. As they form a bounded sequence, and $q(\beta_{t})\in\Delta_{n}$ is also bounded, we can extract a subsequence $({\bar{\beta}}_{t_{s}},q(\beta_{t_{s}}))_{s\in\mathbb{N}}$, with $\lim_{s\rightarrow\infty}t_{s}=\infty$ converging to some limit $({\bar{\beta}}_{\infty},q_{\infty})$. By the Césaro average property, $\frac{1}{t_{s}}\int_{0}^{t_{s}}q(\beta_{s})\mathrm{d}s$ also converges towards $q_{\infty}$. Equation 1 then yields
$\frac{1}{t_{s}}\nabla\phi(\beta_{t_{s}})\underset{s\to\infty}{\longrightarrow}Z^{\top}q_{\infty}$ (2)
@@ 4.3 Assumptions guaranteeing the existence of $\phi_{\infty}$
$\phi_{\infty}({\bar{\beta}})=\lambda\lim_{\eta\to 0}\eta\cdot\varphi^{-1}\left(\phi\left(\frac{{\bar{\beta}}}{\eta}\right)\right)$
@@ 5 Applications and experiments
The hyperbolic entropy potential $\phi^{\mathrm{MD}_{2}}(\beta)=\sum_{i=1}^{d}(\beta_{i}\mathrm{arcsinh}(\beta_{i})-\sqrt{\beta_{i}^{2}+1}-1)$ plays a central role in works considering diagonal linear networks (Woodworth et al., 2020; Pesme and Flammarion, 2023). Applying Theorem 3, we obtain that $\phi_{\infty}\propto\|\cdot\|_{1}$ and we recover the result from Lyu and Li (2020) and Moroshko et al. (2020).
@@
Therefore, by employing various potentials, we can induce different implicit biases, leading to distinct generalisation properties depending on the data distribution.
@@ 6 Conclusion and limitations
In this paper, we offer a comprehensive characterisation of the implicit bias of mirror flow for separable classification problems. This characterisation is framed in terms of the horizon function associated with the mirror descent potential, leveraging the asymptotic geometry induced by the potential.
@@ Appendix A Proofs of properties of the mirror flow in the classification setting
The loss is decreasing. $\frac{\mathrm{d}}{\mathrm{d}t}L(\beta_{t})=-\langle\nabla L(\beta_{t}),\dot{\beta}_{t}\rangle=-\langle\nabla^{2}\phi(\beta_{t})^{-1}\nabla L(\beta_{t}),\nabla L(\beta_{t})\rangle\leq 0$, where the inequality is due to the convexity of the potential $\phi$.
@@ Appendix B Differed proofs on the construction of $\phi_{\infty}$
The sequence of sets $(\bar{S}_{c})$ is contained in the compact ball $B(0,1)$; therefore, Hausdorff convergence is equivalent to Painlevé-Kuratowski convergence [Rockafellar and Wets, 1998, Section 4.C]. Hence, as $(\bar{S}_{c})$ are convex, so is their limit $S_{\infty}$ [Rockafellar and Wets, 1998, Prop 4.15]. It follows that $h_{\infty}$ is convex [Rockafellar and Wets, 1998, Ex 3.50].
Since $S_{\infty}$ is non-degenerate, there exists a radius $r_{0}$ such that $B(0,r_{0})\subset S_{\infty}$, which implies that $h_{\infty}(\beta)$ is finite-valued for every $\beta$.
@@
Consider now a sequence $(\beta_{s})_{s>0}$ satisfying the conditions described in (ii). Since it diverges to infinity and $h$ is coercive, we have $h(\beta_{s})\rightarrow\infty$, and we may assume w.l.o.g that $h(\beta_{s})>0$ for all $s$. We have by definition of sublevel sets $\beta_{s}\in S_{h(\beta_{s})}$, and therefore
$\nabla h(\beta_{s})\in\partial I_{S_{h(\beta_{s})}}(\beta_{s}),$ (8)
which can be derived easily from convexity of $h$ (geometrically, this means that the gradients of $h$ are normal to the sublevel sets). Consider now the normalized levels sets as defined in (4). Denoting
$\bar{\beta}_{s}=\frac{\beta_{s}}{R_{h(\beta_{s}})},$
we have $\bar{\beta}_{s}\in\bar{S}_{h(\beta_{s})}$ and thus by simple rescaling (8) becomes
$\nabla h(\beta_{s})\in\partial I_{\bar{S}_{h(\beta_{s})}}\left(\bar{\beta}_{s}\right).$
Since $\partial I_{\bar{S}_{c}}$ is a cone (the normal cone to $\bar{S}_{c}$), this also holds for any positive multiple of $\nabla h(\beta_{s})$. We deduce that for every $s>0$
$\left(\bar{\beta}_{s},\frac{\nabla h(\beta_{s})}{\|\nabla h(\beta_{s})\|}\right)\in\mathcal{G}(\partial\eta_{h(\beta_{s})}).$
Note that since $\bar{\beta}_{s}$ belongs to the normalized level sets, this sequence is bounded. We can extract a subsequence $(\bar{\beta}_{s_{k}},\frac{\nabla h(\beta_{s_{k}})}{\|\nabla h(\beta_{s_{k}})\|})_{k\geq 0}$ which converges to a limit point $(\hat{\beta},\hat{g})$. By the previous remark on graphical convergence of subdifferentials, we have $(\hat{\beta},\hat{g})\in\mathcal{G}(\partial I_{S_{\infty}})$, i.e.,
@@
We need to prove that $\hat{\beta}$ is not 00. Since $h$ is strictly convex, the level set $\{\phi(\beta)=c\}$ is exactly the boundary of the sublevel set $\{\phi(\beta)\leq c\}$. Therefore, $\beta_{s}$ lies on the boundary of $S_{h(\beta_{s})}$, and hence so does $\bar{\beta}_{s}$ lie on the boundary of $\bar{S}_{h(\beta_{s})}$. Since 00 is in the interior of $S_{\infty}$, it also belongs to the interior of $\bar{S}_{\phi(\beta_{s})}$ for $s$ larger than some $s_{0}$. Then, there exists $r_{0}>0$ such that $B(0,r_{0})\subset\bar{S}_{\phi(\beta_{s})}$ for $s\geq s_{0}$. By definition of boundary, we then have for $s\geq s_{0}$ $\|\bar{\beta}_{s}\|>r_{0}$, which leads to $\|\hat{\beta}\|>0$.
To achieve the desired result; we need to relate $(\hat{\beta},\hat{g})$ to $(\bar{\beta},\bar{g})$. First, notice that by construction we have necessarily $\hat{g}=\bar{g}=\lim_{s\rightarrow\infty}\nabla h(\beta_{s})/\|h(\beta_{s})\|$. Then, note that
$\bar{\beta}=\lim_{k\rightarrow\infty}\frac{\beta_{s_{k}}}{\|\beta_{s_{k}}\|},\quad\hat{\beta}=\lim_{k\rightarrow\infty}\frac{\beta_{s_{k}}}{R_{h(\beta_{s_{k}}})}.$
Taking the norm of the second limit, we have $\|\hat{\beta}\|=\lim_{k\rightarrow\infty}\frac{\|\beta_{s_{k}}\|}{R_{h(\beta_{s_{k}}})}$. Injecting back in the first limit yields
$\bar{\beta}=\lim_{k\rightarrow\infty}\frac{\beta_{s_{k}}}{R_{h(\beta_{s_{k}})}}\frac{R_{h(\beta_{s_{k}})}}{\|\beta_{s_{k}}\|}=\frac{\hat{\beta}}{\|\hat{\beta}\|}.$
@@
$\bar{g}\in\lambda\partial h_{\infty}(\|\hat{\beta}\|\bar{\beta}).$
Finally, $\lambda>0$ since $\|\bar{g}\|=1$, and $\partial h_{\infty}(\|\hat{\beta}\|\bar{\beta})=\partial h_{\infty}(\bar{\beta})$ by positive homogenity of $h_{\infty}$. ∎
@@
$\lim_{t}\frac{1}{t}\nabla\phi(\beta_{t})=Z^{\top}q_{\infty},$ (11)

### v3

@@ 1.1 Informal statement of the main result
Figure 1: Mirror descent is performed using $3$ different potentials on the same toy $2$d dataset. Left: the losses converge to zero. Center: the iterates converge in direction towards $3$ different vectors ${\bar{\beta}}_{\infty}$, the $3$ lines passing through the origin correspond to the associated separating hyperplanes. Right: the limit directions are each proportional to $\mathrm{arg\ min\ }\phi_{\infty}({\bar{\beta}})$ under the constraint $\min_{i}y_{i}\langle x_{i},{\bar{\beta}}\rangle\geq 1$ for their respective $\phi_{\infty}$’s, as predicted by our theory (Theorem 1). The full trajectories are plotted Figure 4 and we refer to Section 5 for more details.
@@ 2 Problem set-up
The first two points of the assumption are commonly used to ensure well-posedness of mirror descent (Bauschke et al., 2017). The third one is necessary in continuous time to ensure the existence and uniqueness over $\mathbb{R}_{\geq 0}$ of a solution to the (MF) differential equation (in particular, we want to avoid the solution “blowing up in finite time”; see Lemma 2 in Appendix A). The coercive gradient assumption is crucial for our main result and we discuss it in more depth in Section 6.
@@ 3 Intuitive construction of the implicit regularisation problem
From now on, we drop the tilde notation and assume that a change of time scale has been done. We want to characterise the directional limit of the diverging iterates $\beta_{t}$. To do so, we study their normalisation $\bar{\beta}_{t}\coloneqq\frac{\beta_{t}}{\|\beta_{t}\|}$. As they form a bounded sequence, and $q(\beta_{t})\in\Delta_{n}$ is also bounded, we can extract a subsequence [^2] More precisely, we let $(t_{n})_{n\in\mathbb{N}}$ be any sequence with $t_{n}\to\infty$ as $n\to\infty$ and then consider the discrete sequence $(\bar{\beta}_{t_{n}},q(\beta_{t_{n}}))_{n\in\mathbb{N}}$. We can then apply the Bolzano–Weierstrass theorem to extract a convergent subsequence. We will show that the limit $(\bar{\beta}_{\infty},q_{\infty})$ is unique and does not depend on the sequence $(t_{n})$, and therefore the continuous-time process $(\bar{\beta}_{t},q(\beta_{t}))_{t\in\mathbb{R}}$ must converge towards it. $({\bar{\beta}}_{t_{s}},q(\beta_{t_{s}}))_{s\in\mathbb{N}}$, with $\lim_{s\rightarrow\infty}t_{s}=\infty$ converging to some limit $({\bar{\beta}}_{\infty},q_{\infty})$. By the Césaro average property, $\frac{1}{t_{s}}\int_{0}^{t_{s}}q(\beta_{s})\mathrm{d}s$ also converges towards $q_{\infty}$. Equation 1 then yields
$\frac{1}{t_{s}}\nabla\phi(\beta_{t_{s}})\underset{s\to\infty}{\longrightarrow}Z^{\top}q_{\infty}.$ (2)
@@ 4.3 Assumptions guaranteeing the existence of $\phi_{\infty}$
$\phi_{\infty}({\bar{\beta}})=\lambda\lim_{\eta\to 0}\eta\cdot\varphi^{-1}\left(\phi\left(\frac{{\bar{\beta}}}{\eta}\right)\right).$
@@ 5 Applications and experiments
The hyperbolic entropy potential: $\phi^{\mathrm{MD}_{2}}(\beta)=\sum_{i=1}^{d}(\beta_{i}\mathrm{arcsinh}(\beta_{i})-\sqrt{\beta_{i}^{2}+1}-1)$ plays a central role in works considering diagonal linear networks (Woodworth et al., 2020; Pesme and Flammarion, 2023). Applying Theorem 3, we obtain that $\phi_{\infty}\propto\|\cdot\|_{1}$ and we recover the result from Lyu and Li (2020) and Moroshko et al. (2020).
@@
Therefore, by employing various potentials, we can induce different implicit biases, leading to distinct generalisation properties depending on the data distribution. The trajectories of the mirror descent iterates are shown and commented in Figure 4.
Figure 4: Mirror flow trajectories on a 2-dimensional dataset for three different potentials (exact same setting as in Figure 1). Left: the iterates diverge to infinity and the directional convergence depends on the choice of potential. Right: the normalised iterates converge towards their respective $\phi_{\infty}$-maximum-margin predictors (illustrated by stars), as predicted by Theorem 2.
@@ 6 Conclusion and limitations
In this paper, we offer a comprehensive characterisation of the implicit bias of mirror flow for separable classification problems. This characterisation is framed in terms of the horizon function associated with the mirror descent potential, leveraging the asymptotic geometry induced by the potential. Note that we did not cover the discrete mirror descent algorithm; we believe the analysis would extend without additional difficulties compared to the continuous counterpart.
Extensions and open problems.
@@ Appendix A Proofs of properties of the mirror flow in the classification setting
The loss is decreasing. $\frac{\mathrm{d}}{\mathrm{d}t}L(\beta_{t})=\langle\nabla L(\beta_{t}),\dot{\beta}_{t}\rangle=-\langle\nabla^{2}\phi(\beta_{t})^{-1}\nabla L(\beta_{t}),\nabla L(\beta_{t})\rangle\leq 0$, where the inequality is due to the convexity of the potential $\phi$.
@@ Appendix B Differed proofs on the construction of $\phi_{\infty}$
The sequence of sets $(\bar{S}_{c})$ is contained in the compact ball $B(0,1)$; therefore, Hausdorff convergence is equivalent to Painlevé-Kuratowski convergence [Rockafellar and Wets, 1998, Section 4.C]. Hence, as $(\bar{S}_{c})$ are convex, so is their limit $S_{\infty}$ [Rockafellar and Wets, 1998, Prop 4.15]. It follows that $\phi_{\infty}$ is convex [Rockafellar and Wets, 1998, Ex 3.50].
Since $S_{\infty}$ is non-degenerate, there exists a radius $r_{0}$ such that $B(0,r_{0})\subset S_{\infty}$, which implies that $\phi_{\infty}(\beta)$ is finite-valued for every $\beta$.
@@
Consider now a sequence $(\beta_{s})_{s>0}$ satisfying the conditions described in (ii). Since it diverges to infinity and $\phi$ is coercive, we have $\phi(\beta_{s})\rightarrow\infty$, and we may assume w.l.o.g that $\phi(\beta_{s})>0$ for all $s$.
We have by definition of sublevel sets that $\beta_{s}\in S_{\phi(\beta_{s})}$, and therefore the gradient $\nabla\phi(\beta_{s})$ belongs to the normal cone of $S_{\phi(\beta_{s})}$ at $\beta_{s}$ (indeed, the gradient is orthogonal to the level sets; see e.g., [Courant and John, 1989, Chapter 1.5]). Since the normal cone of a convex set is the subdifferential of its indicator [Rockafellar, 1970, Section 23], we thus have
$\nabla\phi(\beta_{s})\in\partial I_{S_{\phi(\beta_{s})}}(\beta_{s}),$ (8)
Consider now the normalized levels sets as defined in (4). Denoting
$\bar{\beta}_{s}=\frac{\beta_{s}}{R_{\phi(\beta_{s})}},$
we have $\bar{\beta}_{s}\in\bar{S}_{\phi(\beta_{s})}$ and thus by simple rescaling (8) becomes
$\nabla\phi(\beta_{s})\in\partial I_{\bar{S}_{\phi(\beta_{s})}}\left(\bar{\beta}_{s}\right).$
Since $\partial I_{\bar{S}_{c}}$ is a cone (the normal cone to $\bar{S}_{c}$), this also holds for any positive multiple of $\nabla\phi(\beta_{s})$. We deduce that for every $s>0$
$\left(\bar{\beta}_{s},\frac{\nabla\phi(\beta_{s})}{\|\nabla\phi(\beta_{s})\|}\right)\in\mathcal{G}(\partial\eta_{\phi(\beta_{s})}).$
Note that since $\bar{\beta}_{s}$ belongs to the normalized level sets, this sequence is bounded. We can extract a subsequence $(\bar{\beta}_{s_{k}},\frac{\nabla\phi(\beta_{s_{k}})}{\|\nabla\phi(\beta_{s_{k}})\|})_{k\geq 0}$ which converges to a limit point $(\hat{\beta},\hat{g})$. By the previous remark on graphical convergence of subdifferentials, we have $(\hat{\beta},\hat{g})\in\mathcal{G}(\partial I_{S_{\infty}})$, i.e.,
@@
We need to prove that $\hat{\beta}$ is not 00. Since $\phi$ is strictly convex, the level set $\{\phi(\beta)=c\}$ is exactly the boundary of the sublevel set $\{\phi(\beta)\leq c\}$. Therefore, $\beta_{s}$ lies on the boundary of $S_{\phi(\beta_{s})}$, and hence so does $\bar{\beta}_{s}$ lie on the boundary of $\bar{S}_{\phi(\beta_{s})}$. Since 00 is in the interior of $S_{\infty}$, it also belongs to the interior of $\bar{S}_{\phi(\beta_{s})}$ for $s$ larger than some $s_{0}$. Then, there exists $r_{0}>0$ such that $B(0,r_{0})\subset\bar{S}_{\phi(\beta_{s})}$ for $s\geq s_{0}$. By definition of boundary, we then have for $s\geq s_{0}$ $\|\bar{\beta}_{s}\|>r_{0}$, which leads to $\|\hat{\beta}\|>0$.
To achieve the desired result; we need to relate $(\hat{\beta},\hat{g})$ to $(\bar{\beta},\bar{g})$. First, notice that by construction we have necessarily $\hat{g}=\bar{g}=\lim_{s\rightarrow\infty}\nabla\phi(\beta_{s})/\|\nabla\phi(\beta_{s})\|$. Then, note that
$\bar{\beta}=\lim_{k\rightarrow\infty}\frac{\beta_{s_{k}}}{\|\beta_{s_{k}}\|},\quad\hat{\beta}=\lim_{k\rightarrow\infty}\frac{\beta_{s_{k}}}{R_{\phi(\beta_{s_{k}})}}.$
Taking the norm of the second limit, we have $\|\hat{\beta}\|=\lim_{k\rightarrow\infty}\frac{\|\beta_{s_{k}}\|}{R_{\phi(\beta_{s_{k}})}}$. Injecting back in the first limit yields
$\bar{\beta}=\lim_{k\rightarrow\infty}\frac{\beta_{s_{k}}}{R_{\phi(\beta_{s_{k}})}}\cdot\frac{R_{\phi(\beta_{s_{k}})}}{\|\beta_{s_{k}}\|}=\frac{\hat{\beta}}{\|\hat{\beta}\|}.$
@@
$\bar{g}\in\lambda\partial\phi_{\infty}(\|\hat{\beta}\|\bar{\beta}).$
Finally, $\lambda>0$ since $\|\bar{g}\|=1$, and $\partial\phi_{\infty}(\|\hat{\beta}\|\bar{\beta})=\partial\phi_{\infty}(\bar{\beta})$ by positive homogenity of $\phi_{\infty}$. ∎
@@
$\lim_{t}\frac{1}{t}\nabla\phi(\beta_{t})=Z^{\top}q_{\infty},$